# Policy Improvement using Language Feedback Models

### v3

@@ Abstract
We introduce Language Feedback Models (LFMs) that identify desirable behaviour — actions that help achieve tasks specified in the instruction — for imitation learning in instruction following. To train LFMs, we obtain feedback from Large Language Models (LLMs) on visual trajectories verbalized to language descriptions. First, by using LFMs to identify desirable behaviour to imitate, we improve in task-completion rate over strong behavioural cloning baselines on three distinct language grounding environments (Touchdown, ScienceWorld, and ALFWorld). Second, LFMs outperform using LLMs as experts to directly predict actions, when controlling for the number of LLM output tokens. Third, LFMs generalize to unseen environments, improving task-completion rate by 3.5-12.0% through one round of adaptation. Finally, we modify LFMs to provide human-interpretable feedback without performance loss, allowing human verification of desirable behaviour for imitation learning.
Machine Learning, ICML, Natural Language Processing, instruction following, Language Feedback, Language Grounding, learning feedback model, imitation learning
@@ 1 Introduction
(b) Policy improvement by imitating desirable behaviour identified by a learned feedback model. Given the instruction, we roll out a base policy, then identify productive actions that help achieve tasks specified in the instruction using the trained feedback model. Finally, we update the base policy by imitating the identified desirable behaviour.
@@
(c) Example of desirable behaviour identified in an example environment in ALFWorld, a kitchen instruction following benchmark.
@@
Figure 1: Given an environment and instructions to follow, we assume access to a verbalization procedure that converts observations to language descriptions. Policy improvement using Language Feedback Model involves (a) first training a feedback model, then (b) using it to identify desirable behaviour for policy improvement via imitation learning. We show the feedback model in yellow, other models in purple, and generated intermediate data in green. An example of LFM-identified desirable behaviour is shown in (c).
@@
Sample-efficiency and generalizability are two primary challenges in learning instruction following agents in grounded environments (MacMahon et al., 2006; Kollar et al., 2010; Ahn et al., 2022). First, we want an agent that is sample-efficient: it learns from few demonstrations of how to act according to instructions. Second, we want an agent that is generalizable: it should act successfully in novel environments according to new instructions after training. Reinforcement learning (RL; Sutton & Barto (2018)) and imitation learning (IL; Schaal (1999)Abbeel & Ng (2004)) are two techniques for learning agents for instruction following in grounded environments. These techniques often require large numbers of trials and errors or expensive-to-obtain expert demonstrations. Recent work show that pretrained large language models (LLMs) exhibit sample-efficient learning through prompting and in-context learning for textual (Brown et al., 2020) and grounded problems such as robotic control (Ahn et al., 2022). However, for instruction following in grounded problems, current methods rely on LLMs on-line during inference, which is impractical and expensive.
@@
We develop a sample-efficient and cost-effective technique that uses LLMs to train Language Feedback Models (LFMs) for policy improvement in instruction following. Figure 1 illustrates policy improvement using LFMs. Consider the task of interacting with objects in a kitchen to follow instructions shown in Figure 1(c). First, in Figure 1(a), given a grounded environment and a base policy (i.e. a behaviour cloned policy), we roll out the base policy to collect a small set of trajectories for different instructions. Next, we verbalize observations in the trajectory by describing scenes in language. For each instruction and verbalized trajectory pair, we query an LLM to provide feedback identifying which behaviour in the trajectory is productive to solving the task identified in the instruction (i.e. answer yes or no). For instance, given an instruction “put a clean slice of lettuce in the refridgerator”, GPT-4 (OpenAI, 2023) is able to deduce that key milestones are 1) find the lettuce, 2) slice it 3) wash it in the sink, and 4) put it in the fridge. Consequently, such an LLM is able to identify when an agent is exhibiting desirable behaviour conducive to solving tasks outlined in the instruction, for instance by taking the lettuce to the sink, versus undesirable behaviour, for instance by cooking the lettuce. After collecting LLM feedback, we distill this world knowledge into a small and cost-effective LFM. Finally, in Figure 1(b), given a policy to improve on potentially new environments and instructions, we use the learned LFM to identify desirable actions on-line, then update the policy to imitate these actions. Crucially, this technique is sample-efficient and cost-effective in that it only requires few LLM interactions to collect an off-line dataset during LFM training (i.e. before deployment), as opposed to many LLM interactions on-line during policy improvement (i.e. after deployment).
@@
Our findings are as follows: first, through LFM policy improvement, on three grounded instruction following benchmarks, namely Touchdown (Chen et al., 2019), ScienceWorld (Wang et al., 2022), and ALFWorld (Shridhar et al., 2021b), we observe consistent gains over strong, behaviour cloned base policies. Second, using LLMs as feedback models outperforms using LLMs as expert policies for imitation learning. We compare LFMs against prompting LLMs to directly predict what actions to take, then imitating this LLM-predicted behaviour. On all benchmarks, using LFM feedback outperforms using LLMs as experts for imitation learning, given a fixed allocation of LLM output tokens. This gain is especially pronounced in environments with larger action spaces, such as ScienceWorld, where it is much easier to critique than to generate the correct action. Third, we show that learned feedback models generalize to unseen environments. After training LFMs on training environments, we use them to identify desirable behaviour on test environments, which we then adapt the policy to imitate. A single round of adaptation achieves significant gains (3.5-12.0% task-completion rate) across all environments.
@@
In addition to policy improvement, using LFM feedback offers two advantages over existing techniques such as using LLMs as expert policies for imitation learning. First, LFM improves policies on-line without additional expensive calls to LLMs. Second, LFM can offer human-interpretable feedback when identifying desirable behaviour to imitate. We show in Section 5.4 that LFMs can be easily modified to provide not only desirable behaviour but why they were desirable, thereby allowing humans to inspect and validate imitation data used for policy improvement. Source code for our environments and experiments are available at anonymous.4open.science/r/language_feedback_models. Videos of LFM feedback are available at language-feedback-models.github.io.
@@
Figure 2: An example verbalization for Touchdown. We align Clip image embeddings of panorama patches and language embeddings of common noun-phrases to populate a language template. Appendix A describes this procedure in detail. The blue arrow at the top indicate the agent’s orientation while the green arrows indicate valid directions to proceed in.
@@ 2 Background
In language-grounded instruction following, an agent is given an instruction $x$ specifying the task to achieve in the environment. Each turn, the agent receives a potentially partial observation $o_{t}$, and takes an action $a_{t}$ which causes the environment to transition to a new state. In the example in Figure 1(b), the agent observes a counter with objects such as a toaster, some lettuce, and a knife on top. To follow the instruction “put a clean slice of lettuce in the refridgerator”, an effective agent may choose to grab a piece of lettuce. In the reinforcement learning setting, the environment additionally give the agent a reward after a desirable (positive reward) or undesirable (negative reward) action (Sutton & Barto, 2018). In this work, we consider long-horizon settings with only sparse and delayed task-completion rewards. Consequently, we focus on imitation learning from demonstrations as opposed to reinforcement learning from rewards (Schaal, 1999).
Imitation learning.
In imitation learning for instruction following, we are given an expert policy $\pi^{*}(a|x,o)$ and learn a policy $\pi_{\theta}(a|x,o)$ with parameters $\theta$. We first roll out the policy $\pi_{\theta}$. For each step $o^{(i)}_{t}$ of the rollout $\tau_{i}$, we optimize $\theta$ to imitate the action $a^{(i)}_{t}$ chosen by the expert $\pi^{*}(a|x,o^{(i)}_{t})$ when given the same observations.
$\operatorname*{arg\,min}_{\theta}\operatorname{\mathbb{E}}_{o^{(i)}_{t}\sim\pi_{\theta}}\left[{L}\left(\pi_{\theta}(a|x,o^{(i)}_{t}),a^{(i)}_{t}\right)\right]$ (1)
Here, ${L}$ is a step-wise cross-entropy loss function between the policy’s action distribution and the action chosen by the expert given the same observation:
${L}\left(*\right)=-\sum_{a^{\prime}\in\mathcal{A}}{\mathbbm{1}}\left[a^{\prime}=a^{(i)}_{t}\right]\ln\pi_{\theta}(a=a^{\prime}\mid x,o^{(i)}_{t}).$ (2)
@@
Imitation learning in Eq (1) assumes an expert policy that can be executed on-line to produce expert actions. For instance, given an expert, imitation learning assumes that this expert $\pi^{*}(a|x,o_{t})$ provides corrective actions $a_{t}$ as the policy $\pi(a|x,o_{t})$ runs. In many cases, this is impractical — a human-in-the-loop expert is expensive and inconvenient while an LLM expert is expensive and, as we show in our experiments, inaccurate. Alternatively, in behaviour cloning (BC), we instead collect an offline dataset of expert trajectories from which to clone expert behaviour (Bain & Sammut, 1995; Torabi et al., 2018). BC (or offline imitation learning) only asks the expert to perform the task $N$ times to collect $N$ trajectories $\{\tau_{i}\}_{i=1}^{N}$. Each $\tau_{i}$ consists of $M_{i}$ steps of observations and associated expert actions: $\tau_{i}=[o^{(i)}_{1},a^{(i)}_{1},\ldots,o^{(i)}_{M_{i}},a^{(i)}_{M_{i}}]$ where $a^{(i)}_{t}$ is the action chosen by the expert $\pi^{*}(a|x,o^{(i)}_{t})$ given the observation $o^{(i)}_{t}$. We train policy $\pi_{\theta}$ to imitate the expert action, given the same observation seen by the expert, by minimizing the following objective:
$\operatorname*{arg\,min}_{\theta}\frac{1}{N}\sum_{i}^{N}\frac{1}{M_{i}}\sum_{t}^{M_{i}}{L}\left(\pi_{\theta}(a|x,o^{(i)}_{t}),a^{(i)}_{t}\right).$ (3)
The key distinction between BC and imitation learning is that the former optimizes over trajectories under the expert policy while the latter optimizes over trajectories under the learned policy. Consequently, while BC is offline and easily batchable, it suffers from covariate shift/exposure bias (Ross et al., 2011; Bengio et al., 2015). Like prior work in long-horizon instruction following in grounded environments (Fried et al., 2018; Chen et al., 2019), we use BC to warm-start a strong base policy (Ash & Adams, 2020), which we then improve using imitation learning.
@@ 3 Language Feedback Model
How can we leverage world knowledge in LLMs to make policy learning more sample-efficient and generalizable? In this work, we use LLMs to distill a small and cost-effective Language Feedback Model to identify desirable behaviour from a base policy (Figure 1(a)). We then improve the base policy by imitating this desirable behaviour through batched imitation learning, without need for on-line LLMs (Figure 1(b)). Appendix B provides pseudo-code for the entire procedure for policy improvement using LFMs. A natural question is why not directly use LLMs as experts for action prediction. Section 5.4 shows that the using LLMs to learn feedback models results in higher policy improvement than using LLMs as experts for action prediction. Moreover, LFMs generalize to new environments unseen during training, thereby allowing policy improvement on new environments.
@@ 3.1 Verbalization
To leverage world knowledge in LLMs, we convert raw observations $o$ to language descriptions $v$ using a verbalization procedure $V$. Figure 2 illustrates such a verbalization procedure for Touchdown (Chen et al., 2019), where the agent navigates Google Street View panorama images based on a given natural language instruction. First, we extract all noun-phrases (NPs) from instructions in the dataset and compute their Clip language embedding. Given a visual observation, we compute Clip visual embedding for each image patch, and align it with the top matching NP as deemed by the highest cosine similarity between Clip embeddings. We then combine aligned NPs with agent orientation to formulate an egocentric language description of the scene. This is described in more detail in Appendix A.
@@ 3.2 Learning a feedback model
Given a verbalization procedure $V$, an instruction $x$, an LLM, and a policy $\pi_{\theta}$, we now describe a procedure to use the LLM’s knowledge to improve $\pi_{\theta}$. First, we prompt the LLM to provide feedback on whether a particular action taken by the policy $\pi_{\theta}(a|x,v)$ is productive in achieving the tasks outlined in the instruction $x$. We then improve the policy $\pi_{\theta}$ by updating its parameters to imitate desirable behaviour determined by the LLM. Let :::: denote “such that”. Let ${{\textsc{Llm}}}(x,v,a)$ return ${\mathrm{yes}}$ if and only if the LLM feedback indicates that action $a$ taken in verbalized state $v$ and instruction $x$ is productive. Given a set of instructions $X=\{x_{i}\}_{1}^{N}$, the optimization procedure is then
$\operatorname*{arg\,min}_{\theta}\operatorname{\mathbb{E}}_{v,a^{\prime},x:{{\textsc{Llm}}}(x,v,a^{\prime})={\mathrm{yes}}}{L}\left(\pi_{\theta}\left(a|x,v\right),a^{\prime}\right)$ (4)
where instruction $x$ is sampled from $X$ and the observations $v$ and actions $a^{\prime}$ are sampled from rollouts of the policy $\pi_{\theta}$.
@@
While Eq (4) is a reasonable procedure for using LLM feedback to improve the policy, it requires calling LLMs at each step during policy improvement. This is prohibitively expensive both in terms of query cost, because LLMs capable of giving desirable feedback are expensive to run, and training time, because generating feedback using large LLMs is slow. Instead of using the LLM at each step, we make a modification to the procedure in Eq (4) to collect LLM feedback over long horizons in batch (Colas et al., 2023) in order to train a small and cost-effective language feedback model.
First, for instructions $\{x^{(1)},x^{(2)},\ldots\}$ we roll out the base policy $\pi_{\theta}$ to collect a set of trajectories $\{\tau_{1},\tau_{2},\ldots\}$ consisting of verbalized observations and actions taken: $\tau_{i}=\{v^{(i)}_{1}\pi(x^{(i)},v^{(i)}_{1}),v^{(i)}_{2}\pi(x^{(i)},v^{(i)}_{2}),\ldots\}$. For each $\tau_{i}$, we prompt the LLM for feedback on which steps were productive in achieving the instruction $x^{(i)}$. Table 2’s Lfm row shows an example of requesting feedback from GPT-4 on a rollout in ALFWorld, which is an instruction following benchmark in verbalized 3D kitchens. This LLM feedback is then parsed to identify the precise steps in which the base policy $\pi_{\theta}$ took a productive action towards achieving the goals outlined in the instruction. The set of desirable behaviour is compiled into a dataset $F$. Let $y^{*}={{\textsc{Llm}}}(x,v,a)$ denote the feedback given by the LLM for the instructions $x$, observations $v$, and action $a$. We use the dataset $F=\{x^{(i)},v,a,y^{*}\forall v,a\in\tau_{i}\forall x^{(i)},\tau_{i}\}$ to train a small Language Feedback Model $f$.
$\operatorname*{arg\,min}_{\theta}\sum_{(x,v,a,y^{*})\in F}{L}\left(f_{\theta}\left(y\mid x,v,a\right),y^{*}\right)$ (5)
Here, ${L}$ is the cross-entropy between output distribution of the feedback model $f_{\theta}$ and gold label $y^{*}$ from the LLM.
@@
The naive learning procedure in Eq (4) updates the policy after each step using slow and expensive LLM feedback. Here, we instead update the policy in rounds using fast and cost-effective LFM feedback. In round $k$, we rollout the base policy $\pi^{(k)}$ and use the feedback model $f$ to collect a dataset $D_{k}$ of desirable behaviour. Let $a^{(k)}_{t}$ denote the action chosen by policy $\pi^{(k)}(a\mid x,v_{t})$. Let ${{\textsc{Desirable}}}(x,v,a)=f\left(y={\mathrm{yes}}\mid x,v,a\right)>f\left(y={\mathrm{no}}\mid x,v,a\right)$, returns whether the feedback model predicts that action $a$ is desirable. We have
$D_{k}=\left\{\left(x,v_{t},a^{(k)}_{t}\right)\forall t:{{\textsc{Desirable}}}(x,v_{t},a^{(k)}_{t})\right\}$ (6)
We combine this dataset with previously collected desirable behaviour to update the base policy via imitation learning.
$\theta^{*}=\operatorname*{arg\,min}_{\theta}\sum_{v_{t},a_{t}\in\cup_{i=1}^{k}D_{i}}{L}\left(\pi^{(k)}(a\mid x,v_{t}),a_{t}\right)$ (7)
In the next round, we set the parameters of the base policy $\pi^{(k+1)}$ to be $\theta^{*}$. Should demonstrations be available, we initialize the base policy at $k=1$ to the BC policy, and train on both demonstrations and identified desirable behaviour during subsequent rounds (i.e. $\cup_{i=0}^{k}D_{i}$ where $D_{0}$ is the demonstrations used to train BC).

### v6

@@ Abstract
We introduce Language Feedback Models (LFMs) that identify desirable behaviour — actions that help achieve tasks specified in the instruction — for imitation learning in instruction following. To train LFMs, we obtain feedback from Large Language Models (LLMs) on visual trajectories verbalized to language descriptions. First, by using LFMs to identify desirable behaviour to imitate, we improve in task-completion rate over strong behavioural cloning baselines on three distinct language grounding environments (Touchdown, ScienceWorld, and ALFWorld). Second, imitation learning using LFMs outperform using LLMs as experts to directly predict actions, when controlling for the number of LLM output tokens. Third, LFMs generalize to unseen environments, improving task-completion rate by 3.5-12.0% through one round of adaptation. Finally, we modify LFMs to provide human-interpretable feedback without performance loss, allowing human verification of desirable behaviour for imitation learning.
@@ 1 Introduction
(b) Policy improvement by imitating desirable behaviour identified by a learned feedback model. Given the instruction, we roll out a base policy, then identify productive actions that help achieve tasks specified in the instruction using the trained feedback model. Finally, we update the base policy by imitating productive actions.
@@
(c) Example of desirable behaviour identified in ALFWorld, a kitchen instruction following benchmark.
@@
Figure 1: Given an environment and instructions to follow, we assume a verbalization procedure that converts observations to language descriptions. Policy improvement using Language Feedback Model involves (a) training a feedback model, then (b) using it to identify desirable behaviour for policy improvement via imitation learning. The feedback model is yellow, other models purple, and generated intermediate data green. An example of LFM-identified behaviour is shown in (c).
@@
Sample-efficiency and generalizability are two primary challenges in learning instruction following agents in grounded environments [26, 23, 2]. First, we want an agent that is sample-efficient: it learns from few demonstrations of how to act according to instructions. Second, we want an agent that is generalizable: it should act successfully in novel environments according to new instructions after training. Reinforcement learning (RL; Sutton and Barto [40]) and imitation learning (IL; Schaal [32]Abbeel and Ng [1]) are two techniques for learning agents for instruction following in grounded environments. These techniques often require large numbers of trials and errors or expensive-to-obtain expert demonstrations. Recent work show that pretrained large language models (LLMs) exhibit sample-efficient learning through prompting and in-context learning for textual [10] and grounded problems such as robotic control [2]. However, for instruction following in grounded problems, current methods rely on LLMs on-line during inference, which is impractical and expensive.
@@
We develop a sample-efficient and cost-effective technique that uses LLMs to train Language Feedback Models (LFMs) for policy improvement in instruction following. Figure 1 illustrates policy improvement using LFMs. Consider the task of interacting with objects in a kitchen to follow instructions shown in Figure 1(c). First, in Figure 1(a), given a grounded environment and a base policy (i.e. a behaviour cloned policy), we roll out the base policy to collect a small set of trajectories for different instructions. Next, we verbalize observations in the trajectory by describing scenes in language. For each instruction and verbalized trajectory pair, we query an LLM to provide feedback identifying which behaviour in the trajectory is productive to solving the task identified in the instruction (i.e. answer yes or no). For instance, given an instruction “put a clean slice of lettuce in the refridgerator”, GPT-4 [29] is able to deduce that key milestones are 1) find the lettuce, 2) slice it 3) wash it in the sink, and 4) put it in the fridge. Consequently, such an LLM is able to identify when an agent is exhibiting desirable behaviour conducive to solving tasks outlined in the instruction, for instance by taking the lettuce to the sink, versus undesirable behaviour, for instance by cooking the lettuce. We define desirable behaviour as productive actions that are constructive, task-beneficial, and effective in following the instruction. In other words, taking the action brings the agent closer (in terms of trajectory length) to accomplishing the task specified in the instruction. After collecting LLM feedback, we distill this world knowledge into a small and cost-effective LFM. Finally, in Figure 1(b), given a policy to improve on potentially new environments and instructions, we use the learned LFM to identify desirable actions on-line, then update the policy to imitate these actions. Crucially, this technique is sample-efficient in that it improves policy with no additional human-labeled demonstrations. Furthermore, this technique is cost-effective in that it requires few LLM interactions to collect an off-line dataset during LFM training (i.e. before deployment), as opposed to many LLM interactions on-line during policy improvement (i.e. after deployment).
@@
Our findings are as follows: first, LFM policy improvement achieves consistent gains over strong behaviour cloned base policies on three grounded instruction following benchmarks in Touchdown [12], ScienceWorld [42], and ALFWorld [38]. Second, we compare LFMs against prompting LLMs to directly predict what actions to take, then imitating this LLM-predicted behaviour. On all benchmarks, using LFM feedback outperforms using LLMs as experts for imitation learning, given a fixed allocation of LLM output tokens. This gain is especially pronounced in environments with larger action spaces, such as ScienceWorld, where it is much easier to critique than to generate the correct action. Third, we show that learned feedback models generalize to unseen environments with new tasks and new transition functions. After training LFMs on training environments, we use them to identify desirable behaviour on test environments, which we then imitate to adapt the policy. A single round of adaptation achieves significant gains (3.5-12.0% task-completion) across all environments.
@@
In addition to policy improvement, using LFM feedback offers two advantages over existing techniques such as using LLMs as expert policies for imitation learning. First, LFM improves policies on-line without additional expensive calls to LLMs. Second, LFM can offer human-interpretable feedback when identifying desirable behaviour to imitate. We show in Section 5.4 that LFMs can be easily modified to provide not only desirable behaviour but why they were desirable, thereby allowing humans to inspect and validate imitation data used for policy improvement. Source code for our environments and experiments are available at github.com/vzhong/language_feedback_models. Videos of LFM feedback are available at language-feedback-models.github.io.
@@
Figure 2: An example verbalization for Touchdown. We align Clip image embeddings of panorama patches and language embeddings of common noun-phrases to populate a language template. Appendix D describes this procedure in detail. The blue arrow at the top indicate the agent’s orientation while the green arrows indicate valid directions to proceed in.
@@ 2 Background
In language-grounded instruction following, an agent is given an instruction $x$ specifying the task to achieve in the environment. Each turn, the agent receives a potentially partial observation $o_{t}$, and takes an action $a_{t}$ which causes the environment to transition to a new state. In the Figure 1(c) example, the agent observes a counter with objects such as a toaster, some lettuce, and a knife on top. To follow the instruction “put a clean slice of lettuce in the refridgerator”, an effective agent may choose to grab a piece of lettuce. In the reinforcement learning setting, the environment additionally give the agent a reward after a desirable (positive reward) or undesirable (negative reward) action [40]. In this work, we consider long-horizon settings with only sparse and delayed task-completion rewards. Consequently, we focus on imitation learning from demonstrations as opposed to reinforcement learning from rewards [32].
Online imitation learning.
In online imitation learning for instruction following, we are given an expert policy $\pi^{*}(a|x,o)$ and learn a policy $\pi_{\theta}(a|x,o)$ with parameters $\theta$. We first roll out the policy $\pi_{\theta}$. For each step $o^{(i)}_{t}$ of the rollout $\tau_{i}$, we optimize $\theta$ to imitate the action $a^{(i)}_{t}$ chosen by the expert $\pi^{*}(a|x,o^{(i)}_{t})$ when given the same observations: $\operatorname*{arg\,min}_{\theta}\operatorname{\mathbb{E}}_{\pi^{*}}\left[{L}\left(\pi_{\theta}(a|x,o^{(i)}_{t}),a^{(i)}_{t}\right)\right]$. Here, ${L}$ is step-wise cross-entropy between the policy’s action distribution and the action chosen by the expert given the same observation: ${L}\left(*\right)=-\sum_{a^{\prime}\in\mathcal{A}}{\mathbbm{1}}\left[a^{\prime}=a^{(i)}_{t}\right]\ln\pi_{\theta}(a=a^{\prime}\mid x,o^{(i)}_{t}).$
@@
Online imitation learning assumes an expert policy that can be executed online to produce expert actions. For instance, given an expert, imitation learning assumes that this expert $\pi^{*}(a|x,o_{t})$ provides corrective actions $a_{t}$ as the policy $\pi(a|x,o_{t})$ runs. In many cases, this is impractical — a human-in-the-loop expert is expensive and inconvenient while an LLM expert is expensive and, as we show in our experiments, inaccurate. Alternatively, in behaviour cloning (BC), we instead collect an offline dataset of expert trajectories from which to clone expert behaviour [6, 41]. BC (or offline imitation learning) only asks the expert to perform the task $N$ times to collect $N$ trajectories $\{\tau_{i}\}_{i=1}^{N}$. Each $\tau_{i}$ consists of $M_{i}$ steps of observations and associated expert actions: $\tau_{i}=[o^{(i)}_{1},a^{(i)}_{1},\ldots,o^{(i)}_{M_{i}},a^{(i)}_{M_{i}}]$ where $a^{(i)}_{t}$ is the action chosen by the expert $\pi^{*}(a|x,o^{(i)}_{t})$ given the observation $o^{(i)}_{t}$. We train policy $\pi_{\theta}$ to imitate the expert action, given the same observation seen by the expert, by minimizing the following objective: $\operatorname*{arg\,min}_{\theta}\frac{1}{N}\sum_{i}^{N}\frac{1}{M_{i}}\sum_{t}^{M_{i}}{L}\left(\pi_{\theta}(a|x,o^{(i)}_{t}),a^{(i)}_{t}\right).$ The key distinction between BC and imitation learning is that the former optimizes over trajectories under the expert policy while the latter optimizes over trajectories under the learned policy. Consequently, while BC is offline and easily batchable, it suffers from covariate shift/exposure bias [31, 7]. Like prior work in long-horizon instruction following in grounded environments [18, 12], we use BC to warm-start a strong base policy [4], which we then improve using imitation learning.
@@ 3 Language Feedback Model
How can we leverage world knowledge in LLMs to make policy learning more sample-efficient and generalizable? In this work, we use LLMs to distill a small and cost-effective Language Feedback Model to identify desirable behaviour from a base policy (Figure 1(a)). We then improve the base policy by imitating this desirable behaviour through batched imitation learning, without need for on-line LLMs (Figure 1(b)). Appendix E provides pseudo-code for the entire procedure for policy improvement using LFMs. A natural question is why not directly use LLMs as experts for action prediction. Section 5.4 shows that the using LLMs to learn feedback models results in higher policy improvement than using LLMs as experts for action prediction. Moreover, LFMs generalize to new environments unseen during training, thereby allowing policy improvement on new environments.
@@ 3.1 Verbalization
To leverage world knowledge in LLMs, we convert raw observations $o$ to language descriptions $v$ using a verbalization procedure $V$. Figure 2 illustrates verbalization for Touchdown [12], where the agent navigates Google Street View panorama images based on a given natural language instruction. First, we extract all noun-phrases (NPs) from instructions in the dataset and compute their Clip language embedding. Given a visual observation, we compute Clip visual embedding for each image patch, and align it to the NP with the highest cosine similarity between Clip embeddings. We then combine aligned NPs with agent orientation to formulate an egocentric language description of the scene. This is described in more detail in Appendix D.
@@ 3.2 Learning a feedback model
Given a verbalization procedure $V$, an instruction $x$, an LLM, and a policy $\pi_{\theta}$, we now describe a procedure to use the LLM’s knowledge to improve $\pi_{\theta}$. First, we prompt the LLM to provide feedback on whether a particular action taken by the policy $\pi_{\theta}(a|x,v)$ is productive in achieving the tasks outlined in the instruction $x$. We then improve the policy $\pi_{\theta}$ by updating its parameters to imitate desirable behaviour determined by the LLM. Let :::: denote “such that”. Let ${{\textsc{Llm}}}(x,v,a)$ return ${\mathrm{yes}}$ if and only if the LLM feedback indicates that action $a$ taken in verbalized state $v$ and instruction $x$ is productive. Given a set of instructions $X=\{x_{i}\}_{1}^{N}$, the optimization procedure is then $\operatorname*{arg\,min}_{\theta}\operatorname{\mathbb{E}}_{v,a^{\prime},x:{{\textsc{Llm}}}(x,v,a^{\prime})={\mathrm{yes}}}{L}\left(\pi_{\theta}\left(a|x,v\right),a^{\prime}\right)$. Here, instruction $x$ is sampled from $X$. Observations $v$ and actions $a^{\prime}$ are sampled from rollouts of the policy $\pi_{\theta}$.
@@
While the previously described naive learning is a reasonable procedure for using LLM feedback to improve the policy, it requires calling LLMs at each step during policy improvement. This is prohibitively expensive both in terms of query cost, because LLMs capable of giving desirable feedback are expensive to run, and training time, because generating feedback using large LLMs is slow. Instead of using the LLM at each step, we make a modification to collect LLM feedback over long horizons in batch [14] in order to train a small and cost-effective language feedback model.
First, for instructions $\{x^{(1)},x^{(2)},\ldots\}$ we roll out the base policy $\pi_{\theta}$ to collect a set of trajectories $\{\tau_{1},\tau_{2},\ldots\}$ consisting of verbalized observations and actions taken: $\tau_{i}=\{v^{(i)}_{1}\pi(x^{(i)},v^{(i)}_{1}),v^{(i)}_{2}\pi(x^{(i)},v^{(i)}_{2}),\ldots\}$. For each $\tau_{i}$, we prompt the LLM for feedback on which steps were productive in achieving the instruction $x^{(i)}$. Table 2’s Lfm row shows an example of requesting feedback from GPT-4 on a rollout in ALFWorld, which is an instruction following benchmark in verbalized 3D kitchens. This LLM feedback is then parsed to identify the precise steps in which the base policy $\pi_{\theta}$ took a productive action towards achieving the goals outlined in the instruction. The set of desirable behaviour is compiled into a dataset $F$. Let $y^{*}={{\textsc{Llm}}}(x,v,a)$ denote the feedback given by the LLM for the instructions $x$, observations $v$, and action $a$. We use the dataset $F=\{x^{(i)},v,a,y^{*}\forall v,a\in\tau_{i}\forall x^{(i)},\tau_{i}\}$ to train a small Language Feedback Model $f$:
$\operatorname*{arg\,min}_{\theta}\sum_{(x,v,a,y^{*})\in F}{L}\left(f_{\theta}\left(y\mid x,v,a\right),y^{*}\right).$ (1)
Here, ${L}$ is the cross-entropy between the feedback model output $f_{\theta}$ and gold label $y^{*}$ from the LLM.
@@
The naive learning procedure in Eq (3.2) updates the policy after each step using slow and expensive LLM feedback. Here, we instead update the policy in rounds using fast and cost-effective LFM feedback. In round $k$, we rollout the base policy $\pi^{(k)}$ and use the feedback model $f$ to collect a dataset $D_{k}$ of desirable behaviour. Let $a^{(k)}_{t}$ denote the action chosen by policy $\pi^{(k)}(a\mid x,v_{t})$. Let ${{\textsc{Desirable}}}(x,v,a)=f\left(y={\mathrm{yes}}\mid x,v,a\right)>f\left(y={\mathrm{no}}\mid x,v,a\right)$ return whether the feedback model predicts that action $a$ is desirable. In the $k$th round, we collect the dataset of desirable behaviour $D_{k}=\left\{\left(x,v_{t},a^{(k)}_{t}\right)\forall t:{{\textsc{Desirable}}}(x,v_{t},a^{(k)}_{t})\right\}$, which we combine with previously collected behaviour to update the policy via imitation learning:
$\theta^{*}=\operatorname*{arg\,min}_{\theta}\sum_{v_{t},a_{t}\in\cup_{i=1}^{k}D_{i}}{L}\left(\pi^{(k)}(a\mid x,v_{t}),a_{t}\right).$ (2)
In the next round, we set the base policy $\pi^{(k+1)}$ parameters to $\theta^{*}$. Should demonstrations be available, we initialize the base policy at $k=1$ to the BC policy, and train on both demonstrations and identified desirable behaviour during subsequent rounds (i.e. $\cup_{i=0}^{k}D_{i}$ where $D_{0}$ are demos used to train BC).